# Parameter-Efficient Fine-Tuning for Large Models: A Comprehensive Survey

### v5

@@ VI-A System design for PEFT
In this section, we begin by providing a concise overview of cloud-based PEFT systems. Following this, we present the corresponding metrics employed for evaluating the system performance. Additionally, we present three prospective utilization scenarios to illustrate the challenges in system design.
@@ VI-A1 Centralized PEFT Query Serving
Cloud providers have recently introduced a range of LLM services aimed at providing user applications through application programming interfaces (APIs) [239, 240]. These APIs facilitate the seamless integration of many ML functionalities into applications. After receiving one query for one specific downstream task through API, the cloud-based server processes the query with one featured LLM model. Under this scenario, the proposed cloud solution for handling multiple PEFT queries involves storing only a single copy of the LLM and multiple PEFT modules. This single copy maintains multiple branches of PEFT modules, each associated with different PEFT queries. The case study of a state-of-the-art system can be found in Section VI-C. Figure 10 (b) illustrates the computation pattern for multi-query PEFT inference, wherein packed PEFT queries are scheduled and executed according to their deadlines and current system conditions.
VI-A2 Serving Metrics
To evaluate the system performance of centralized PEFT query serving, we propose a set of evaluation metrics.
@@ VI-A2 Serving Metrics
VI-A3 Distributed System for PEFT
Nevertheless, in the contemporary LLM model, personalized tasks are not fully supported with pre-trained models, consequently, extra fine-tuning is required to be executed with the methodologies mentioned in the previous sections. However, a big concern is raised when we consider giving the datasets to cloud providers since these datasets are personalized.
For this concern, DLoRA [241] presents a distributed PEFT framework. During the PEFT process, the backbone LLM is executed in the cloud servers while the PEFT modules are trained entirely within the user devices. DLoRA scheme is depicted in Figure 10(a).
Figure 10: (a) Distributed-based system computation pattern; (b) centralized PEFT Query inference
VI-A4 Distributed Metrics
To assess the efficacy of the proposed method, we establish a set of evaluative metrics. For this analysis, and without loss of generality, we adopt language models as the basis for our metric definitions.
@@ VI-A4 Distributed Metrics
  • Compute cost: The compute cost during forward and backward propagation operations on edge devices.
@@
VI-A5 Multi-PEFT Training
Different from multiple-PEFT serving, tuning with multiple customized PEFTs always involves different backbone LLMs. When contemplating LLM usage across various downstream tasks, pre-trained models typically exhibit subpar performance. A prevalent approach to adapt LLM to diverse tasks involves crafting fine-tuned PEFTs. However, simultaneously tuning multiple PEFTs can pose considerable challenges. Challenges like how to manage memory gradient and model weights storage, and how to design an efficient kernel for batching PEFT training remain unsolved. PEFTs will be categorized based on their PEFT algorithms and backbone LLM models. The design challenge involves how to consolidate multiple PEFTs with the same LLM backbone and multiple different LLM backbones simultaneously.
VI-B Case study: Offsite-Tuning
To resolve these two issues, Offsite-Tuning [242] proposes a privacy-preserving and efficient transfer learning framework that enables foundational models to adapt to downstream tasks without the need to access the complete model weights. The key insight of Offsite-Tuning is the cloud provider sends an adapter and an emulator to the data proprietor. Then, with the assistance of the emulator, the data proprietor fine-tunes the adapter. The fine-tuned adapter is then sent back to the cloud side, which integrates it into the complete model, creating a fine-tuned foundational model for downstream users.
Offsite-Tuning safeguards the privacy of data proprietors since they do not need to share their training data directly. It also protects the foundational model owners, as the complete model weights are not shared, and the emulator provided is lossy, with significantly degraded performance. Compared to existing fine-tuning methods that require access to the full model weights, Offsite-Tuning is more resource-efficient because it allows for fine-tuning through a compressed emulator without needing the complete model.
VI-C Case Study: PetS
@@ VI-D Parallel PEFT Training Frameworks
Design Challenges
Unlike the PetS system, which aims to accommodate flexible multi-PEFT algorithms, S-LoRA [244] and Punica [245] focus solely on facilitating multiple-LoRA blocks for various tasks. Designing multiple PEFT training systems presents key challenges in two main aspects:
@@ VII-A Simplify hyperparameter tuning
The effectiveness of PEFT is often sensitive to its hyperparameters, such as the bottleneck dimension of the adapter, the rank of LoRA, and the arrangement of various additive PEFT layers. Manually tuning these hyperparameters will cost lots of efforts. Therefore, future efforts could focus on developing methods that are less dependent on manual tuning of these parameters, or automatically find the optimal configuration settings. Several studies [76, 77, 78, 91, 92, 93] have started to address this issue, but there’s a need for more simple and efficient solutions optimizing these hyperparameters.
@@ VII-C Enhance training efficiency
The presumed parameter efficiency of PEFT does not always consistent with computational and memory savings during training. Given that trainable parameters are intertwined within the pre-trained model’s architecture, computing and storing activations and gradients for the full model often become necessary during fine-tuning. This oversight calls for a rethinking of what constitutes efficiency. As outlined in Section IV, potential solutions lie in the integration of model compression techniques such as pruning and quantization, alongside innovations specifically designed to optimize memory during PEFT tuning [249]. Further research into enhancing the computational efficiency of PEFT methodologies is imperative.
@@ VII-G PEFT with model compression
Model compression is one of the most effective ways to make LLM executable on resource-limited devices. Yet, the impact of model compression techniques on the performance of PEFT algorithms running on hardware remains another systemic challenge. Common compression techniques such as quantization and pruning necessitate dedicated hardware platforms to expedite the process, and building such hardware platform for compressed models is yet another direction for future research.

### v7

@@ VI-A System design for PEFT
In this section, we begin by providing a concise overview of cloud-based PEFT systems and analyzing the design challenges. These include the efficient handling of numerous task-specific queries via centralized PEFT query servicing, the resolution of privacy and data transmission issues through distributed PEFT training, and the complexities associated with concurrent multi-PEFT training processes. Centralized systems are required to process a substantial volume of queries with minimal latency and maximal throughput. Distributed training frameworks must address privacy concerns and the computational inefficiencies that arise from data exchanges between users and cloud services. Furthermore, multi-PEFT training necessitates the optimization of memory utilization, the management of simultaneous model training, and the formulation of system architectures capable of supporting multi-tenant workloads effectively. These challenges underscore the imperative for innovative approaches to improve scalability, safeguard privacy, and optimize resource allocation in PEFT system architectures. Following this, we present the corresponding metrics employed for evaluating the system performance. Furthermore, we delve into three prospective utilization scenarios to illustrate the challenges in system design.
@@ VI-A1 Centralized PEFT Query Serving
Cloud providers have recently introduced a range of LLM services aimed at providing user applications through application programming interfaces (APIs) [246, 247]. These APIs facilitate the seamless integration of many machine-learning functionalities into applications. When receiving one query for one specific downstream task through API, the cloud-based server processes the query with one featured LLM model. Under this scenario, the importance of PEFT becomes apparent. Cloud providers store only a single copy of the LLM and multiple PEFT modules featuring different downstream tasks. This setup allows the LLM to maintain various branches of PEFT modules, each linked to specific API queries, i.e., PEFT queries.
Centralized PEFT query serving solutions address scenarios where multiple PEFT queries arrive in quick succession. A case study of one state-of-the-art system for this purpose is discussed in Section VI-B. Figure 10 (b) illustrates the computation pattern for multi-query PEFT inference, wherein packed PEFT queries are scheduled and executed according to their deadlines and current system conditions.
VI-A2 Distributed PEFT Training
In most cases, personalized tasks are not fully supported with pre-trained models, consequently, extra fine-tuning is required to be executed with the methodologies mentioned in the previous sections. However, significant concerns arise when considering the transfer of datasets to cloud providers, given the issues related to data privacy, copyright, proprietary information, and the complexities and inefficiencies involved in data transmission. Section VI-C gives two approaches that address this concern.
Figure 10: (a) Distributed-based system computation pattern; (b) centralized PEFT Query inference.
VI-A3 Multi-PEFT Training
Different from multiple-PEFT serving, tuning with multiple customized PEFTs always involves different backbone LLMs. Therefore, simultaneously tuning multiple PEFTs can pose considerable challenges. Challenges like how to manage memory gradient and model weights storage, and how to design an efficient kernel for batching PEFT training remain unsolved. PEFTs will be categorized based on their PEFT algorithms and backbone LLM models. The design challenge involves how to consolidate multiple PEFTs with the same LLM backbone and multiple different LLM backbones simultaneously. We present case studies related to this topic in Section VI-D.
VI-A4 Evaluation Metrics
For the proposed evaluation metrics, without loss of generality, we adopt large language models as the basis for our metric definitions.
To evaluate the system performance of PEFT serving systems, we propose a set of evaluation metrics:
@@ VI-A4 Evaluation Metrics
To assess the efficacy of PEFT training systems, we also establish a set of evaluative metrics:
@@
  • Compute cost: The compute cost during forward and backward propagation operations on cloud servers and edge devices.
@@
VI-B Centralized PEFT Serving Frameworks
@@
The other example design is DLoRA [249], which introduces a system that improves the efficiency of serving low-rank adaptation (LoRA) models for large language models (LLMs) by dynamically managing the merging and unmerging of LoRA adapters and the migration of requests across worker replicas. This dynamic orchestration addresses the challenges of high memory footprints, low GPU utilization, and load imbalance caused by variable input and output lengths in traditional LLM serving systems. dLoRA’s novel approaches, including a credit-based batching algorithm and a request-adapter co-migration algorithm, significantly enhance throughput.
VI-C Distributed PEFT Training Frameworks
To resolve these two issues, DLoRA [250] presents a distributed PEFT framework. During the PEFT process, the backbone LLM is executed in the cloud servers while the PEFT modules are trained entirely within the user devices. DLoRA scheme is depicted in Figure 10(a).
Similarly, Offsite-Tuning [251] presents a privacy-preserving and efficient transfer learning framework that enables foundational models to adapt to downstream tasks without the need to access the complete model weights. The key insight of Offsite-Tuning is the cloud provider sends an adapter and an emulator to the data proprietor. Then, with the assistance of the emulator, the data proprietor fine-tunes the adapter. The fine-tuned adapter is then sent back to the cloud side, which integrates it into the complete model, creating a fine-tuned foundational model for downstream users. Offsite-Tuning safeguards the privacy of data proprietors since they do not need to share their training data directly. It also protects the foundational model owners, as the complete model weights are not shared, and the emulator provided is lossy, with significantly degraded performance. Compared to existing fine-tuning methods that require access to the full model weights, Offsite-Tuning is more resource-efficient because it allows for fine-tuning through a compressed emulator without needing the complete model.
@@ VI-D Parallel PEFT Training Frameworks
Unlike the PEFT query serving system, which aims to accommodate flexible multi-PEFT algorithms, Punica [252] focuses solely on facilitating multiple-LoRA blocks for various tasks. Designing multiple PEFT training systems presents key challenges in two main aspects:
@@ VII-A Simplify hyperparameter tuning
The effectiveness of PEFT is often sensitive to its hyperparameters, such as the bottleneck dimension of the adapter, the rank of LoRA, and the arrangement of various additive PEFT layers. Manually tuning these hyperparameters will cost lots of effort. Therefore, future efforts could focus on developing methods that are less dependent on manual tuning of these parameters, or automatically find the optimal configuration settings. Several studies [82, 83, 84, 98, 99, 100] have started to address this issue, but there’s a need for more simple and efficient solutions optimizing these hyperparameters.
@@ VII-C Enhance training efficiency
The presumed parameter efficiency of PEFT is not always consistent with computational and memory savings during training. Given that trainable parameters are intertwined within the pre-trained model’s architecture, computing and storing activations and gradients for the full model often become necessary during fine-tuning. This oversight calls for a rethinking of what constitutes efficiency. As outlined in Section IV, potential solutions lie in the integration of model compression techniques such as pruning and quantization, alongside innovations specifically designed to optimize memory during PEFT tuning [256]. Further research into enhancing the computational efficiency of PEFT methodologies is imperative.
@@ VII-G PEFT with model compression
Model compression is one of the most effective ways to make LLM executable on resource-limited devices. Yet, the impact of model compression techniques on the performance of PEFT algorithms running on hardware remains another systemic challenge. Common compression techniques such as quantization and pruning necessitate dedicated hardware platforms to expedite the process, and building such hardware platforms for compressed models is yet another direction for future research.